# Shape-specific fluctuations of an active colloidal interface

### v2

@@ V.2 Orientational fluctuations
The orientational fluctuations are displayed in Fig. 3(A) (and inset). There are two regimes displayed: (i) Ballistic regime. This corresponds to deterministic orientational changes as the C-shape forms. (ii) Steady-state regime. $W_{\theta}$ is distinguished from $W_{h}$ is that there is no $N$ dependent $t^{*}$ during which the fluctuations relax. A rough estimate of this from Fig. 3(A) would be $t_{\theta}^{*}\sim 50\tau_{f}$, thus of the same order as $t_{h}^{*}(N)$. Further, $l_{c}$ spans the system size at all times as displayed in Fig. 3(A) inset.
@@
We further report that $W_{\theta}^{SS}$ scales negatively with the chain length. We may thus call this a smoothness exponent. This form of orientational stabilization with the system size, as opposed to conventional “roughness” seen in other growing interfaces. The source of emergence for such a smoothening exponent is thus the increased orientational rigidity along the chain, which increases with chain length [25]. The angular distribution along the C-shape becomes increasingly uniform, thereby making the chain stiffer and more polarized along the x-axis. The RMS deviations from the the mean orientation thus decreases with system size (a larger fraction of the chain is polarized along the x-axis). This is in contrast to the positional steady-state fluctuations, where the mean $h(y)$ varies with chain length, and fluctuations about this exhibits roughness.
@@ V.3 With finite rotational noise
Addition of rotational noise may indeed vary the possible DSS attained from simulation. We do not report them here; but simply mention that $\tau_{r}\approx 10$ sets a threshold above which the C-shape is the attractor. In this section we further investigate whether the aforementioned findings carry over to the case of finite $\tau_{r}$. Indeed we find that $\tau_{r}\gtrsim 20$ an identical FV scaling for $W_{h}$ is obtained, whereas $W_{\theta}$ only approximates the scaling law and collapse (15) - See Fig. 4. For $W_{h}$ one sees a crossover at $\tau_{r}$ for the angular fluctuations between a diffusive $\beta_{\theta}\approx 0.5$ and a deterministic $\beta_{\theta}\approx 1$. The effect of randomness prohibits a systems-spanning $l_{c}$. Nevertheless, a smoothness exponent with an identical scaling can be obtained.
For $\tau_{r}\lesssim 10$ there is no longer a propelling C-shape, and indeed other DSS are observed. The reason for this is that the C-shape DSS is sensitive to long-wavelength fluctuations in this regime, such that any fluctuations on the length scale $\frac{b^{2}D_{r}}{v_{s}}$ is greater than typical monomer lengths. It is to be noted that for equilibrated systems $\tau_{r}=\frac{8\pi\eta b^{3}}{k_{b}T}$, with $\tau_{t}=\frac{6\pi\eta b^{2}}{k_{b}T}$ ($\eta$ the viscosity of the medium). Thus the effect of rotational noise and hence potential loss of the C-shape DSS may set in for smaller $b$. This trade-off and its potential experimental relevance for the system considered here is discussed further in the Discussion.
@@ V.4 Locally flat regimes
In the study of fluctuations of circular interfaces, the fluctuations are typically measured about a locally flat region along the growing front [36, 37, 9]; and the roughness scalings thus correspond to those of a locally flat region. We repeat the same procedure here, and vary $N^{\prime}$ under the constraint of $N^{\prime}<<N$. In this section, $\frac{N^{\prime}}{N}<0.18$ such that the interface is effectively flat. The results are presented in Fig. 5. Probing the locally flat regime, we find that the height fluctuations in the early-time ($\beta^{(1)}$) and steady-state regimes, respectively, can be collapsed onto a FV-type scaling with growth exponent $\beta_{h}^{\prime}\approx 0.25$, and roughness exponent $\alpha_{h}^{\prime}\approx 1$. This would imply a dynamic exponent $z_{h}^{\prime}=4$. This combination of exponents is reminiscent of interfaces with conservation laws in two dimensions  [38, 17, 41, 42]. However, it is to be noted that this implies essentially neglecting the super-ballistic regime, where the data does not collapse. The $W_{\theta}$ scaling in this regime is presented in Appendix .6.
@@ VI Summary and Discussions
Roto-translational coupling has been well-studied for their various implications on dynamical and collective behaviour in active matter physics. The effect of this coupling on scaling laws have been established for the mean-squared displacement - i.e enhanced or anomalous diffusion - in various models (see for instance [33, 43, 44]). Here, we instead present the effect of roto-translation-induced global topological change, and the resultant novel scaling laws for interface fluctuations. Remarkably, we find that the height fluctuations show an FV data collapse, with a new set of dynamic and roughness exponents; with the former rationalized via a simple linearised analysis of roto-translational coupling. The orientational fluctuations, in addition, display novel scaling behaviour, notably smoothening with system size, which we attribute to the enhanced orientational rigidity of longer chains.
Some of our observed exponents are rationalizable via comparison with those of well-studied models. The roughness exponent 0.9~0.9 accompanying the height fluctuations is higher than that of non-conserving mass models ($\sim 0.5$) and lower than conserving mass models ($\sim 1.5$). These two correspond to the limit of “monomeric units” exchanging with the bulk (e.g. evaporation) and being distributed across the surface respectively. The latter limit for instance corresponds to the dynamic exponent we report in the locally flat regime (with monomeric redistribution but discounting the C-shape). We thus see why the C-shape roughness lies in between these two values - the C-shape does distribute monomeric displacements across the body (thus less suppression of large scale fluctuations), but does so whilst maintaining a fixed bending rigidity. The growth exponent here being super-ballistic is solely a consequence of topological change due to roto-translational coupling; a simple picture as to how roto-translational coupling was give rise to super-ballistic growth is provided in Appendix .3, where an exact solution to a linearized approximate dynamics is presented (a flat chain with lowest-order roto-translational coupling). For the orientational sector, the smoothness can be appreciated by noting that as the system size increases, a proportionately larger fraction of the monomers are polarized along the x-axis, thus orientational fluctuations at larger length scales get more suppressed. The ballistic growth exponent on the other hand arises due to the deterministic effect of orientational changes that are disproportionately on the chain edges.
@@
It is of further relevance to note where these results - the existence of growth and roughness exponents that are unique to this particular topology - fit in the broader study of active interfaces/membranes. Firstly, it is noted that non-equilibrium extensions to passive (growing) interfaces models have been theoretically proposed and new universality classes proposed. For instance, studying the capillary fluctuations on a non-equilibrium field theory, it has been shown that (in 2D) $\alpha=1/2$ and $z=3$, whilst additional non-linear terms decreases these values slightly [45, 46, 23]. These interfaces arise from underlying scalar fields derived from conservation laws whose dynamics violate symmetries found in equilibrium. A parallel set of active membrane models are those where a Monge-like membrane is coupled to (the concentration field) a membrane-spanning protein pump that exerts a mechanical force on the membrane upon exchanging material with it (chemical coupling), with solutions to these exhibiting travelling waves [47, 48, 49]. A recent set of works study a variant of this where the protein dynamics are inclusions on the height field which hop along the membrane with an additional time scale [19, 50, 51]. For these sets of “active membrane” models, the height correlations are characterized by $S(q)=\langle h(\mathbf{q})h(\mathbf{-q})\rangle\sim 1/q^{2}$ universally, the result for a Monge (passive) interface with Gaussian fluctuations [52]. This gives rise to the standard $\alpha=1/2$ roughening. For orientational correlations, though these have not been studied elsewhere to our knowledge, perhaps the closest set of works pertain to those studying tangential angular correlations on a liquid–liquid interface undergoing phase separation, where the tangential angular correlations show a smoothening at large scales [21, 53]. It is evident that the physics of these aforementioned systems are starkly different from that of the autophoretic C-shape chain; we nevertheless mention them here for the sake of completion.
@@
It remains an open challenge to verify these exponents in an experimental system. To observe significant positional fluctuations in experimental realizations, if one equates damping forces to thermal forces, $\eta bv_{s}\sim k_{b}T/b$ ($\eta$ the viscosity of the medium and $v_{s}$ the self-propulsion $\approx 10\mu ms^{-1}$) (e.g. for droplet experiments [25, 54, 55]), one obtains $b\approx 10^{-7}m$, thus around 10 times smaller than typical chemically interacting colloidal systems. On the other hand, other phoretic colloidal systems (typically electrophoretic and bubble propulsion mechanisms) have reported propulsion speeds of $v_{s}\sim 1\mu ms^{-1}$ [56, 57], where the thermal fluctuations will be appreciable. Constructing extended propelling objects with a well-defined propelling interface out of these units remain an open challenge. The smaller colloids of $b\sim 10^{-7}m$ deserve a specific consideration. These correspond to so-called “nanomotor” systems, sub-micron (mostly metallic) colloids whose synthesis routes are well established [58, 59] and have been shown to self-propel via various mechanisms [60, 61, 62, 63]. Noting that $\tau_{r}\propto b^{3}$, we find that with $b\sim 10^{-7}m$, $\tau_{r}\sim 10^{-2}s$, which we note is smaller than the regime required for C-shape formation to take place (c.f. Section V.3). Thus, it may be feasible to only observe the orientational growth and smoothening with an appreciable $b\sim 10^{-6}m$, where $\tau_{r}\sim 10^{1}s$, which supports the C-shape. At this scale it is also important to note that motion is no longer restricted to be in two dimensions, and continuum theories of phoretic/fluid interactions may in addition break down [64, 65]. In this context it is of interest to note recent works that have reported how phase change in 8CB liquid crystal emulsions (micron-sized) can be induced via external temperature changes [54]. In the nematic/isotropic phases, the individual colloids no longer merely deterministically self-propel but in addition have a fluctuating propulsion direction. This would correspond to finite $\tau_{r}$ in our model without sacrificing monomeric size. Creating extended objects out of such systems, and/or realizing a robust C-shape with positional fluctuations thus remain an open problem. Finally, though extended colloidal chains propelled via external (as opposed to autophoretic) mechanisms have widely been synthesized [66, 67, 68, 69, 26, 27], neither have them have reported well defined shapes where interfacial height deviations and fluctuations may be readily computed - i.e. neither of these reported shapes that propel perpendicular to the body axis [70, 29]. It is noted that these systems suffer from the same translational noise suppression and noise scaling issues mentioned above.
Appendix/SI
@@
$\Delta\mathbf{r}_{ij}$ $=\mathbf{r}_{i}(t)-\mathbf{r}_{j}(t^{\prime}),\qquad\mathcal{D}=D_{c}|t-t^{\prime}|.$ (19)
@@
We present the results for $W_{h}$ and $W_{\theta}$ in Fig. 6 (A). We find that the scalings of $W_{h}$ are reproduced. With the micellar trails, when one starts such a system with arbitrary initial conditions (e.g. random), the trails deposited via (19) give rise to a forward-backward symmetry breaking in the orientational response to the chemicals. Thus, the system breaks the initial symmetry imposed on it and further universally picks up the C-shape. In our system, we choose sufficiently random initial conditions such that this effect is seen; we then compute $\Delta h$ once the chain has a well-defined interface.
For this section only one system size, $N=64$, was studied, averaged over $10$ realizations. Note that the dynamical evolution via (19) requires at each time step a sum over all previous time steps. Thus, even for this system size (on a desktop computer with processor frequency of 2.40GHz and 16 cores), one realization takes $3.77\times 10^{5}$ seconds. Thus, a rigorous system-size determination of roughening/smoothening is prohibitive. With only one curve for $W$, obtaining $\alpha$ and $\beta$ via (12) (14) requires self-consistency. For $W_{h}$, we fix $\alpha\approx 0.9$; in this setting the best fit obtained was $\beta_{h}=1.41\pm 0.01$. We have further checked that super-ballistic $\beta_{h}$ is always obtained if the coarsening is in between evaporative and mass conserving limits (this is solely a consequence of the C-shape topology and its fixed bending rigidity, irrespective of dynamical details; see arguments in Section VI).
Further, for $W_{\theta}$, we find that fixing $\beta_{\theta}=1$ gives $\alpha_{\theta}\approx 0.2$. Thus, not a smoothening exponent, but nevertheless much less rough than the standard growth models. If one attempts to fit a negative $\alpha_{\theta}$ instead, one instead finds super-ballistic orientational growth (e.g. $\alpha_{\theta}\approx-0.1$ gives $\beta_{\theta}\approx 4$). Super-ballisticity could also be anticipated given that the time evolution contains higher order temporal derivatives, specifically for monomers near the edges (e.g. $i=9,50,55$ in Fig.6(C) versus $i=311$ of Fig.3(D)), thus any fluctuations about the mean will deviate from a simple ballistic scaling. On the other hand, the steady-state angular profile is no longer fully polarized along the propulsion direction (compare Fig.6(B) versus Fig.3(C)), thus $\alpha_{\theta}$ is expected to be higher than in the case without trails.
@@
$\frac{\partial h_{i}}{\partial t}=v_{s}+\mu\Big(|h_{i+1}-h_{i}-2b|-|h_{i}-h_{i-1}-2b|\Big)+\xi_{i,t}$ (20)
@@
$\frac{\partial h}{\partial t}=v_{s}+\tilde{\mu}\frac{\partial^{2}h}{\partial y^{2}}+\xi$ (21)
@@
$P(W^{2},t)=\int_{-i\infty}^{i\infty}\frac{dy}{2\pi i}e^{xy}\prod_{n=1}^{\infty}\frac{-ys^{2}_{n0}e^{-n^{2}\tau}/(1+ya_{n})}{\langle W^{2}\rangle_{SS}\left(1+ya_{n}\right)},$ (22)
@@
$x=\frac{W^{2}}{W^{2}_{SS}},~\tau_{p}=\frac{8\pi^{2}\nu_{h}t}{L^{2}},~s_{n0}^{2}=\frac{2|a_{n}(0)|^{2}}{W^{2}_{SS}}.$ (23)
@@
$P(W^{2},t)=\sum_{m=1}^{\infty}\frac{1}{a_{m}}\frac{\exp\left[\frac{-x}{a_{m}}\right]}{\langle W^{2}\rangle_{SS}}\prod_{n=1,\neq m}^{\infty}\frac{a_{m}}{a_{m}-a_{n}}.$ (24)
@@
The early time approx of (21) corresponds to setting $\Lambda\rightarrow\infty$. Note that one can either $\tau<<\tau_{f}$ probe or $\chi_{r}=0$ globally; we pick the latter. The analytical distribution of (24) can then be compared to this limit; we find that the distribution is reproduced - black line and blue scatter of Fig. 7(A). We can thus suitably compare this with the case of $\Lambda$ finite, restricting to $\tau<<\tau_{f}$. Here, we find that our propelled interface displays deviations from this passive distribution; this is displayed in Fig. 7(B) - yellow scatter. For both cases, we plot alongside the best fit analytical distribution in dashed black, and we further find that the averaged least squared error of the fit is substantially larger in the case of $\Lambda$ finite; we thus conclude that the early time non-equilibrium distribution deviates from the passive EW case. This marginal deviation at early times arises again via the aforementioned roto-translational coupling; this thus again constitute a non-equilibrium signature that appears via deterministic (but small) orientational interactions. This deviation in the height distribution is analogous to that seen in so-called “anomalous diffusion” [73, 74], where microscopic colloidal particles in specific systems display a mean-squared displacement exponent of $1$, but the probability distribution of displacements has been found to follow the Laplace (instead of Gaussian) distribution [74, 75]. The exact analytical distribution of the chemically interacting chain thus remains an open future problem.
@@
In the main text we have reported the growth exponent of the C-shape to be $\beta_{h}\sim 1.7$ (super-ballistic). Although rationalizing this for a curved surface is beyond exact analytical calculation in this work; we study instead a flat interface, via a the solution to the linearised approximate dynamics.
Let us study the simplest case of roto-translational coupling, in particular where (small) angular dynamics set in; in this setting the full set of equations are solvable. Consider the $i$th positional dynamics from (2). For small $\theta_{i}$, we can write this as
$\dot{x}_{i}=v_{s}\left(1-\frac{\theta_{i}^{2}}{2}\right)+\mu F^{b}_{x_{i}}+\xi_{t,i}$ (25)
@@
$\dot{\theta}_{i}=\chi_{r}\left(1-\frac{\theta_{i}^{2}}{2}\right)S_{i}+\xi_{r,i}$ (26)
where $S_{i}=\frac{1}{4b^{3}}\sum_{j}{\frac{1}{(i-j)^{2}}\Big[2\Theta(i-j)-1\Big]}$ for a completely flat chain. Here, $\Theta(x)$ is the step function. This simplification allows us to write a linearized (26), without accounting for chemical currents along the $x$ direction. We note that fluctuations along the body axis ($y$ direction) mean that $S_{i}$ is random quantity at each time step. The deterministic solution to this is
$\theta_{i}=\sqrt{2}\frac{1-\exp[-\chi_{r}S_{i}t]}{\exp[-\chi_{r}S_{i}t]+1}$ (27)
We next probe the early time regime of $t<<\tau_{f}$; this corresponds to the approximation of a flat surface (c.f. Fig. 1(d)). We then plug (27) into (25). We then use the small argument expansion for $\cosh(z)\approx 1+\frac{z^{2}}{2}$, with $z=\frac{\chi_{r}S_{i}t}{2}$. This gives us
$\dot{x}_{i}=F_{i}(t)\approx v_{s}\Big(1-\frac{b^{6}}{2\tau_{f}^{2}}S_{i}^{2}t^{2}\Big)+F^{b}_{x,i}+\xi_{t,i}$ (28)
@@
$\langle X\rangle=v_{s}t-\frac{b^{6}}{6\tau_{f}^{2}}\langle S^{2}\rangle t^{3}$ (29)

### v3

@@ V.2 Orientational fluctuations
The orientational fluctuations are displayed in Fig. 3(A) (and inset). There are two regimes displayed: (i) Ballistic regime. This corresponds to deterministic orientational changes as the C-shape forms. (ii) Steady-state regime. $W_{\theta}$ is distinguished from $W_{h}$ in that there is no $N$ dependent $t^{*}$ during which the fluctuations relax. A rough estimate of this from Fig. 3(A) would be $t_{\theta}^{*}\sim 50\tau_{f}$, thus of the same order as $t_{h}^{*}(N)$. Further, $l_{c}$ spans the system size at all times as displayed in Fig. 3(A) inset.
@@
We further report that $W_{\theta}^{SS}$ scales negatively with the chain length [Fig. 3(B)]. We may thus call this a smoothness exponent. This form of orientational stabilization with the system size, as opposed to conventional “roughness”, and implies a suppression of long-ranged orientational fluctuations at larger system sizes. The source of emergence for such a smoothening exponent is the increased orientational rigidity along the chain, which increases with chain length [25]. The angular distribution along the C-shape becomes increasingly uniform, thereby making the chain stiffer and more polarized along the x-axis. The RMS deviations from the the mean orientation thus decreases with system size (a larger fraction of the chain is polarized along the x-axis). This is in contrast to the positional steady-state fluctuations, where the mean $h(y)$ varies with chain length, and fluctuations about this exhibits roughness.
@@ V.3 With finite rotational noise
The exponents discussed so far correspond to the case $D_{r}=0$ (equivalently $\tau_{r}\rightarrow\infty$) in (1). Incorporating finite rotational noise $D_{r}$ may indeed change the possible DSS obtained from simulations. We do not report those here; but simply mention that $\tau_{r}\approx 10$ sets a threshold above which the C-shape is the attractor. In this section, we further investigate whether the aforementioned findings carry over to the case of finite $\tau_{r}$. Indeed, we find that $\tau_{r}\gtrsim 20$ an identical FV scaling for $W_{h}$ is obtained, whereas $W_{\theta}$ only approximates the scaling law and collapse (12) - See Fig. 4. The effect of randomness prohibits a system-spanning $l_{c}$. Nevertheless, a smoothness exponent with an identical scaling can be obtained.
For $\tau_{r}\lesssim 10$ there is no longer a propelling C-shape, and indeed other DSSs are observed. The reason for this is that the C-shape DSS is sensitive to long-wavelength fluctuations in this regime, such that any fluctuations on the length scale $\frac{b^{2}D_{r}}{v_{s}}$ is greater than typical monomer lengths. It is to be noted that for equilibrated systems $\tau_{r}=\frac{8\pi\eta b^{3}}{k_{b}T}$, with $\tau_{t}=\frac{6\pi\eta b^{2}}{k_{b}T}$ ($\eta$ is the viscosity of the medium). Thus the effect of rotational noise and hence potential loss of the C-shape DSS may set in for smaller $b$. This trade-off and its potential experimental relevance for the system considered here is discussed further in Sec. VI.
@@ V.4 Locally flat regimes
In the study of fluctuations of circular interfaces, the fluctuations are typically measured about a locally flat region along the growing front [36, 37, 9]; and the roughness scalings thus correspond to those of a locally flat region. We repeat the same procedure here, and vary $N^{\prime}$ under the constraint of $N^{\prime}<<N$. In this section, $\frac{N^{\prime}}{N}<0.18$ such that the interface is effectively flat. The results are presented in Fig. 5. Probing the locally flat regime, we find that the height fluctuations in the early-time ($\beta^{(1)}$) and steady-state regimes, respectively, can be collapsed onto a FV-type scaling with growth exponent $\beta_{h}^{\prime}\approx 0.25$, and roughness exponent $\alpha_{h}^{\prime}\approx 1$. This would imply a dynamic exponent $z_{h}^{\prime}=4$. This combination of exponents, interestingly, is reminiscent of interfaces with conservation laws in two dimensions  [38, 17, 41, 42]. However, it is to be noted that this implies essentially neglecting the super-ballistic regime, where the data does not collapse. The $W_{\theta}$ scaling in this regime is presented in Appendix .6.
@@ VI Summary and Discussions
Roto-translational coupling has been well-studied for their various implications on dynamical and collective behaviour in active matter physics. The effect of this coupling on scaling laws has been investigated for the mean-squared displacement - i.e enhanced or anomalous diffusion - in various models (see for instance [33, 43, 44]). Here, we instead present the effect of roto-translation-induced global topological change, and the resultant novel scaling laws for interface fluctuations. Remarkably, we find that the height fluctuations show an FV data collapse, with a new set of dynamic and roughness exponents; with the growth exponent rationalized via a simple linearised analysis of our model equations presented in Sec. II. The orientational fluctuations, in addition, display novel scaling behaviour, notably smoothening with system size, which we attribute to the enhanced orientational rigidity of longer chains.
Some of our observed exponents are rationalizable via comparison with those of well-studied models. The roughness exponent 0.9~0.9 accompanying the height fluctuations is higher than that in non-conserved models of total height ($\sim 0.5$[34, 35] and lower than that in conserved models ($\sim 1.5$) [17]. These two correspond to the limit of “monomeric units” exchanging with the bulk (e.g. evaporation) and being distributed across the surface respectively. The latter limit for instance corresponds to the dynamic exponent we report in the locally flat regime (with monomeric redistribution but discounting the C-shape). We thus see why the C-shape roughness lies in between these two values - the C-shape does distribute monomeric displacements across the body (thus less suppression of large scale fluctuations), but does so whilst maintaining a fixed bending rigidity. The growth exponent here being super-ballistic is solely a consequence of topological change due to roto-translational coupling; a simple picture as to how roto-translational coupling could give rise to super-ballistic growth is provided in Appendix .3, where an exact solution to a linearized approximate dynamics is presented (a flat chain with lowest-order roto-translational coupling). For the orientational sector, the smoothness can be appreciated by noting that as the system size increases, a proportionately larger fraction of the monomers are polarized along the x-axis, thus orientational fluctuations at larger length scales get more suppressed. The ballistic growth exponent on the other hand arises due to the deterministic effect of orientational changes that take place near the chain edges.
@@
It is of further relevance to note where these results - the existence of growth and roughness exponents that are unique to this particular topology - fit in the broader study of active interfaces/membranes. Firstly, it is noted that non-equilibrium extensions to passive interface models have been theoretically investigated with propositions of new universality classes. For instance, by studying capillary fluctuations using a non-equilibrium field theory, it has been shown that (in 2D) $\alpha=1/2$ and $z=3$, whilst additional non-linear terms decrease these values slightly [45, 46, 23]. These interfaces arise from underlying scalar fields derived from conservation laws whose dynamics violate symmetries found in equilibrium. A parallel set of active membrane models have been studied where a fluctuating membrane is coupled to a membrane-spanning protein pump (the concentration field) that exerts a mechanical force on the membrane upon exchanging material with it (chemical coupling), leading to traveling waves [47, 48, 49]. A recent set of works study a variant of this in which the protein dynamics is modeled as inclusions in the height field that move along the membrane with an additional time scale [19, 50, 51]. For these sets of “active membrane” models, the height correlations are characterized by $S(q)=\langle h(\mathbf{q})h(\mathbf{-q})\rangle\sim 1/q^{2}$ universally, the result for a Monge (passive) interface with Gaussian fluctuations [52]. This gives rise to the standard $\alpha=1/2$ roughening. For orientational correlations, though these have not been studied elsewhere to our knowledge, perhaps the closest set of works pertain to those studying tangential angular correlations on a liquid–liquid interface undergoing phase separation, where the tangential angular correlations show a smoothening at large scales [21, 53]. It is evident that the physics of these aforementioned systems are starkly different from that of the autophoretic C-shape chain; we nevertheless mention them here for the sake of completion.
@@
It remains an open challenge to verify our predicted exponents in an experimental system. To observe significant positional fluctuations in experimental realizations, if one equates damping forces to thermal forces, $\eta bv_{s}\sim k_{b}T/b$ ($\eta$ the viscosity of the medium and $v_{s}$ the self-propulsion $\approx 10\mu ms^{-1}$) (e.g. for droplet experiments [25, 54, 55]), one obtains $b\approx 10^{-7}m$, thus around 10 times smaller than typical chemically interacting colloidal systems. On the other hand, other phoretic colloidal systems (typically electrophoretic and bubble propulsion mechanisms) have reported propulsion speeds of $v_{s}\sim 1\mu ms^{-1}$ [56, 57], where the thermal fluctuations will be appreciable. Constructing extended propelling objects with a well-defined propelling interface out of these units remain an open challenge. The smaller colloids of $b\sim 10^{-7}m$ deserve a specific consideration. These correspond to so-called “nanomotor” systems, sub-micron (mostly metallic) colloids whose synthesis routes are well established [58, 59] and have been shown to self-propel via various mechanisms [60, 61, 62, 63]. Noting that $\tau_{r}\propto b^{3}$, we find that with $b\sim 10^{-7}m$, $\tau_{r}\sim 10^{-2}s$, which we note is smaller than the regime required for C-shape formation to take place (c.f. Section V.3). Thus, it may be feasible to only observe the orientational growth and smoothening with an appreciable $b\sim 10^{-6}m$, where $\tau_{r}\sim 10^{1}s$, which supports the C-shape. At this scale it is also important to note that motion is no longer restricted to be in two dimensions, and continuum theories of phoretic/fluid interactions may in addition break down [64, 65]. In this context it is of interest to note recent works that have reported how phase change in 8CB liquid crystal emulsions (micron-sized) can be induced via external temperature changes [54]. In the nematic/isotropic phases, the individual colloids no longer merely deterministically self-propel but in addition have a fluctuating propulsion direction. This would correspond to finite $\tau_{r}$ in our model without sacrificing monomeric size. Creating extended objects out of such systems, and/or realizing a robust C-shape with positional fluctuations thus remain an open problem. Finally, though extended colloidal chains propelled via external (as opposed to autophoretic) mechanisms have widely been synthesized [66, 67, 68, 69, 26, 27], none of them have reported well-defined shapes where interfacial height deviations and fluctuations may be readily computed - i.e. none of these reported shapes propel perpendicular to the body axis [70, 29]. It is noted that these systems suffer from the same translational noise suppression and noise scaling issues mentioned above.
VII Appendix
@@
$\Delta\mathbf{r}_{ij}$ $=\mathbf{r}_{i}(t)-\mathbf{r}_{j}(t^{\prime}),\qquad\mathcal{D}=D_{c}|t-t^{\prime}|.$ (A1)
@@
We present the results for $W_{h}$ and $W_{\theta}$ in Fig. 6 (A). We observe that the super-ballistic growth regime of $W_{h}$ is reproduced. With the micellar trails, when one starts such a system with arbitrary initial conditions (e.g. random), the trails deposited via (A1) give rise to a forward-backward symmetry breaking in the orientational response to the chemicals. Thus, the system breaks the initial symmetry imposed on it and further universally picks up the C-shape. In our system, we choose sufficiently random initial conditions such that this effect is seen; we then compute $\Delta h$ once the chain has a well-defined interface.
For this section only one system size, $N=64$, was studied, averaged over $10$ realizations. Note that the dynamical evolution via (A1) requires at each time step a sum over all previous time steps. Thus, even for this system size (on a desktop computer with processor frequency of 2.40GHz and 16 cores), one realization takes $3.77\times 10^{5}$ seconds. Thus, a rigorous system-size determination of roughening/smoothening is prohibitive. With only one curve for $W$, obtaining $\alpha$ and $\beta$ via (9) (11) requires self-consistency. For $W_{h}$, we fix $\alpha\approx 0.9$; in this setting the best fit obtained was $\beta_{h}=1.41\pm 0.01$. We have further checked that super-ballistic $\beta_{h}$ is always obtained if the coarsening is in between evaporative and height conserving limits (this is solely a consequence of the C-shape topology and its fixed bending rigidity, irrespective of dynamical details; see arguments in Section VI).
Further, for $W_{\theta}$, we find that fixing $\beta_{\theta}=1$ gives $\alpha_{\theta}\approx 0.2$ – thus, not a smoothening exponent, but nevertheless much less rough than the standard growth models. If one attempts to fit a negative $\alpha_{\theta}$ instead, one instead finds super-ballistic orientational growth (e.g. $\alpha_{\theta}\approx-0.1$ gives $\beta_{\theta}\approx 4$). Super-ballisticity could also be anticipated given that the time evolution contains higher order temporal derivatives, specifically for monomers near the edges (e.g. $i=9,50,55$ in Fig.6(C) versus $i=311$ of Fig.3(D)), thus any fluctuations about the mean will deviate from a simple ballistic scaling. On the other hand, the steady-state angular profile is no longer fully polarized along the propulsion direction (compare Fig.6(B) versus Fig.3(C)), thus $\alpha_{\theta}$ is expected to be higher than in the case without trails.
@@
$\frac{\partial h_{i}}{\partial t}=v_{s}+\mu\Big(|h_{i+1}-h_{i}-2b|-|h_{i}-h_{i-1}-2b|\Big)+\xi_{i,t}$ (A2)
@@
$\frac{\partial h}{\partial t}=v_{s}+\tilde{\mu}\frac{\partial^{2}h}{\partial y^{2}}+\xi$ (A3)
@@
$P(W^{2},t)=\int_{-i\infty}^{i\infty}\frac{dy}{2\pi i}e^{xy}\prod_{n=1}^{\infty}\frac{-ys^{2}_{n0}e^{-n^{2}\tau}/(1+ya_{n})}{\langle W^{2}\rangle_{SS}\left(1+ya_{n}\right)},$ (A4)
@@
$x=\frac{W^{2}}{W^{2}_{SS}},~\tau_{p}=\frac{8\pi^{2}\nu_{h}t}{L^{2}},~s_{n0}^{2}=\frac{2|a_{n}(0)|^{2}}{W^{2}_{SS}}.$ (A5)
@@
$P(W^{2},t)=\sum_{m=1}^{\infty}\frac{1}{a_{m}}\frac{\exp\left[\frac{-x}{a_{m}}\right]}{\langle W^{2}\rangle_{SS}}\prod_{n=1,\neq m}^{\infty}\frac{a_{m}}{a_{m}-a_{n}}.$ (A6)
@@
The early time approximation of (A3) corresponds to setting $\Lambda\rightarrow\infty$. Note that one can either probe $\tau<<\tau_{f}$ or $\chi_{r}=0$ globally; we pick the latter. The analytical distribution of (A6) can then be compared to this limit; we find that the distribution is reproduced - black line and blue scatter of Fig. 7(A). We can thus suitably compare this with the case of $\Lambda$ finite, restricting to $\tau<<\tau_{f}$. Here, we find that our propelled interface displays deviations from this passive distribution; this is displayed in Fig. 7(B) - yellow scatter. For both cases, we plot alongside the best fit analytical distribution in dashed black, and we further find that the averaged least squared error of the fit is substantially larger in the case of $\Lambda$ finite; we thus conclude that the early time non-equilibrium distribution deviates from the passive EW case. This marginal deviation at early times arises again via the aforementioned roto-translational coupling; this thus again constitutes a non-equilibrium signature that appears via deterministic (but small) orientational interactions. This deviation in the height distribution is analogous to that seen in so-called “anomalous diffusion” [73, 74], where microscopic colloidal particles in specific systems display a mean-squared displacement exponent of $1$, but the probability distribution of displacements has been found to follow the Laplace (instead of Gaussian) distribution [74, 75]. The exact analytical distribution of the chemically interacting chain thus remains an open future problem.
@@
In the main text we have reported the growth exponent of the C-shape to be $\beta_{h}\sim 1.7$ (super-ballistic). Although rationalizing this for a curved surface is beyond exact analytical calculation in this work; we study instead a flat interface, via a solution to the approximate linearised dynamics.
Let us study the simplest case of roto-translational coupling, in particular, where (small) angular dynamics set in; in this setting the full set of equations are solvable. Consider the $i$th monomers’ positional dynamics from (1). For small $\theta_{i}$, we can write this as
$\dot{x}_{i}=v_{s}\left(1-\frac{\theta_{i}^{2}}{2}\right)+\mu F^{b}_{x_{i}}+\xi_{t,i}$ (A7)
@@
$\dot{\theta}_{i}=\chi_{r}\left(1-\frac{\theta_{i}^{2}}{2}\right)S_{i}+\xi_{r,i}$ (A8)
where $S_{i}=\frac{1}{4b^{3}}\sum_{j}{\frac{1}{(i-j)^{2}}\Big[2\Theta(i-j)-1\Big]}$ for a completely flat chain. Here, $\Theta(x)$ is the step function. This simplification allows us to write a linearized (A8), without accounting for chemical currents along the $x$ direction. We note that fluctuations along the body axis ($y$ direction) mean that $S_{i}$ is random quantity at each time step. The deterministic solution to this is
$\theta_{i}=\sqrt{2}\frac{1-\exp[-\chi_{r}S_{i}t]}{\exp[-\chi_{r}S_{i}t]+1}$ (A9)
We next probe the early time regime of $t<<\tau_{f}$; this corresponds to the approximation of a flat surface (c.f. Fig. 1(d)). We then plug (A9) into (A7). We then use the small argument expansion for $\cosh(z)\approx 1+\frac{z^{2}}{2}$, with $z=\frac{\chi_{r}S_{i}t}{2}$. This gives us
$\dot{x}_{i}=F_{i}(t)\approx v_{s}\Big(1-\frac{b^{6}}{2\tau_{f}^{2}}S_{i}^{2}t^{2}\Big)+F^{b}_{x,i}+\xi_{t,i}$ (A10)
@@
$\langle X\rangle=v_{s}t-\frac{b^{6}}{6\tau_{f}^{2}}\langle S^{2}\rangle t^{3}$ (A11)